OpenStudy (anonymous):

Given x^2+4xy+y^2=-12, find the equations of all horizontal tangent lines. I already found that y'= -x/2+y

OpenStudy (across):

Implicit differentiation yields\[x^2+4xy+y^2=-12,\]\[2x+4y+4x\frac{dy}{dx}+2y\frac{dy}{dx}=0,\]\[2x+4y+\frac{dy}{dx}\left(4x+2y\right)=0,\]\[\frac{dy}{dx}\left(4x+2y\right)=-2x-4y,\]\[\frac{dy}{dx}=-\frac{x+2y}{2x+y},\]\[-\frac{x}{2}-y=0.\]

OpenStudy (anonymous):

oh wow i forgot to use chain rule thanks

OpenStudy (anonymous):

wait so how do i find the horizontal tangent lines?

OpenStudy (mr.math):

The tangent line is horizontal when of a function \(y=f(x)\) is horizontal if \(\frac{dy}{dx}=0\), and that's what across did when she set up \(\frac{dy}{dx}=-\frac{x+2y}{y+2x}=0\). This expression is zero if, and only if the numerator is zero, thus the line \(x+2y=0\) represents all horizontal tangent lines to \(x^2+4xy+y^2=-12\).

OpenStudy (mr.math):

Sorry, The tangent line of a function \(y=f(x)\) is horizontal if ...